### v1

# XFormParser: A Simple and Effective Multimodal Multilingual Semi-structured Form Parser

Abstract.
In the domain of document AI, semi-structured form parsing plays a crucial role and is employed across a broad spectrum of academic and industrial applications, including structured data management, knowledge retrieval, and content generation. This task encompasses a complex form of data mining that leverages techniques from key information extraction (KIE), dealing with inputs that range from plain text to intricate modal data comprising images and structural layouts. The advent of pre-trained multimodal models significantly empowers document AI frameworks to adeptly extract pivotal information from form documents presented in diverse formats such as PDFs and images. Nonetheless, the endeavor of form parsing is still encumbered by notable challenges like subpar capabilities in multi-lingual parsing and diminished recall in contexts rich in text and visuals. In this work, we introduce a simple but effective Multimodal and Multilingual semi-structured FORM PARSER (XFormParser), which anchored on a comprehensive pre-trained language model and innovatively amalgamates semantic entity recognition (SER) and relation extraction (RE) into a unified framework, enhanced by a novel staged warm-up training approach that employs soft labels to significantly refine form parsing accuracy without amplifying inference overhead. Furthermore, we have developed a groundbreaking benchmark dataset, named InDFormBench, catering specifically to the parsing requirements of multilingual forms in various industrial contexts. Through rigorous testing on established multilingual benchmarks and InDFormBench, XFormParser has demonstrated its unparalleled efficacy, notably surpassing the state-of-the-art (SOTA) models in RE tasks within language-specific setups by achieving an F1 score improvement of up to 1.79%. Our framework exhibits exceptionally improved performance across tasks in both multi-language and zero-shot contexts when compared to existing SOTA benchmarks. The robustness and reliability of XFormParser are thoroughly validated through comprehensive evaluations on the InDFormBench benchmark. [^1] The codes, datasets, and pre-trained models will be released.
Semi-structured form, Information extraction, Multi-modal, multilingual knowledge, Pre-trained language model
conference: ; ; submissionid: 286
1. Introduction
Document AI is the technology of automatically reading, understanding, and analyzing business documents. It is widely applied in commercial, governmental, and educational sectors and is crucial to departmental efficiency and productivity. A key task of Document AI is to parse and extract form information from scanned documents. Form parsing is essentially an entity relation mining task that connects the Named Entity Recognition (NER) task (Li et al., 2020) and Key Information Extraction (KIE) task (Cui et al., 2021; Yu et al., 2021; Hong et al., 2022). Due to the diversity of the layout and the format, poor quality of scanned document images, and complexity of template structures, representing and understanding the unstructured information in documents using generic rules becomes a highly challenging task.
Figure 1. An illustration of named entity recognition for unstructured forms.
Unlike traditional NER tasks (Schuster et al., 2013; Esser et al., 2012) which only deal with textual information, and traditional pattern recognition tasks (Dengel and Klein, 2002; Medvet et al., 2011, 2011), form parsing requires consideration of multiple modalities. Apart from textual information, the position and layout of text blocks also play a crucial role in semantic interpretation. Previously, rules and workflows for each type of document often needed to be hard, requiring updates with changes in specific formats or when dealing with multiple formats. Current researchers believe that effective language models for form parsing need to comprehend target entities and adapt to different document formats in contexts involving multiple modalities. To address these issues, advanced Form Document AI models and algorithms are expected to automatically classify, extract, and structure information from business documents, minimizing manual intervention.
Mainstream document intelligence methods typically involve using deep neural networks to model elements in documents from the perspectives of computer vision (He et al., 2016), natural language processing (Devlin et al., 2018), or Multimodal fusion (Atrey et al., 2010). Form parsing algorithms based on deep learning initially involved detecting and classifying specific regions of document images. Then, it utilized Optical Character Recognition (OCR) models (Li et al., 2023) to extract text information. This process is called unstructured to semi-structured, and the resulting form is called semi-structured. Subsequently, different categories of form blocks and their corresponding text content were routed to dedicated information extraction modules, constructing a pipeline to process entire-page documents(Xu et al., 2020b; Wang et al., 2022).
With the development of pre-training techniques (Zoph et al., 2020) and multimodal fusion technologies, researchers have developed a plethora of semi-structured form parsing algorithms based on Transformer (Vaswani et al., 2017). (Yu et al., 2021; Zhang et al., 2020) proposed a graph convolutional neural network information extraction model that fuses textual and visual information for form parsing in specific business scenarios. (Xu et al., 2020a) introduced the LayoutLM pre-trained model, which combined language models and Transformer, expanding BERT (Devlin et al., 2018) architecture by incorporating layout information to consider spatial relationships between tokens and textual content connections, resulting in outstanding performance in form parsing tasks. Furthermore, LayoutLMv2 (Xu et al., 2020b) and LayoutLMv3 (Huang et al., 2022) enhance the integration of multimodal information in the pretraining stage. They also add pretraining strategies for text-image alignment and text-image matching, incorporating word token alignment targets. LayoutLMv3 is the first multimodal model in Document AI that does not rely on pre-trained CNN (He et al., 2016) or Faster R-CNN (Girshick, 2015) backbones to extract visual features, further promoting the training efficiency and generalizability of the model. The XFormParser(Chen et al., 2023) model proposes a Global Structure Knowledge-guided relation extraction framework. The loop uses the ”generate-capture-merge” operation so that entity representation and global structural knowledge reinforce each other.
Although these models have made significant progress in the field of Semi-structured form parsing, as research and applications have become more widespread, most existing methods still face three limitations: 1) The accuracy of key information extraction from complex, multilingual form images remains insufficient; 2) Multimodal models dominated by visual modalities still lag behind text-dominated multimodal models in form parsing tasks in rich-text and long-text scenarios; 3) General Multi-modal Large Language Models (MLLM), like GPT4 (Huang et al., 2023) and their training data are not completely open-source, making it impossible to directly access advanced pre-trained weights and tracks. Therefore, based on the self-supervised pre-training model, it is crucial to further mine the representation methods of entities and the relationships between entities in the background of multi-modal and multi-lingual knowledge and to study more effective fine-tuning paradigms in complex scenes for understanding the mechanism of table information extraction and effectively applying it to the industrial field.
To address these issues, we propose a semi-structured form parser with multimodal and multilingual knowledge, named XFormParser. For input data from semi-structured forms, XFormParser utilizes a text-centric pre-trained model to generate vectors containing text information, visual information, and spatial positional information. Subsequently, these vectors are fed into the downstream joint network to complete two tasks: Semantic Entity Recognition (SER) and Relation Extraction (RE), namely NER and KIE tasks for semi-structured form, to realize form parsing. We use the multilingual document understanding pre-trained model LayoutXLM (Xu et al., 2022) as the backbone of XFormParser. The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM (Chung et al., 2014; Sun et al., 2022) and Biaffine (Nguyen and Verspoor, 2019). In addition, we further build InDFormBench, a Chinese and English multi-scenario form parsing dataset for industrial applications, based on public benchmarks. Training the model on this dataset helps XFormParser learn semi-structured forms from the real world and achieve new SOTA performance.
The contributions of this paper are summarized as follows:
  • We propose XFormParser, an efficient model for multi-language semi-structured form parsing based on a pre-trained Transformer. It integrates two tasks: SER and RE, along with the joint loss function and training method of soft labels warm-up in stages. XFormParser effectively enhances form parsing performance without additional inference resources and overhead.
  • We construct InDFormBench, a cross-scenario form parsing benchmark dataset in both Chinese and English. It contains 500 form images collected from eight major industrial application scenarios, along with corresponding annotation information in JSON format semi-automatically generated and rigorously validated using GPT4.
  • Through experiments and analysis, we confirm that XFormParser achieves the F1 score of at most 1.79% over the SOTA model for RE tasks in Language-specific scenarios. XFormParser achieves significantly better results than SOTA for both dual-task in Multi-Language and Zero-shot scenarios. We also verify the effectiveness and reliability of XFormParser through experiments on InDFormBench.
2. Related Work
2.1. Single Modality-Based Extraction Methods.
In recent years, various document AI tasks based on natural language understanding (NLU) or computer vision (CV) have made great progress (alm; Yang et al., 2022c, a; Mo et al., 2024; Cheng et al., 2024; Yang et al., 2024; Wang et al., 2023; Ro et al., 2020). CloudScan (Palm et al., 2017) attempts to use LSTM to extract information from invoice-type form documents, on this basis, Sage et al. (Sage et al., 2019) added contextual semantic and spatial information for information extraction. Nguyen et al. (Nguyen and Verspoor, 2019), obtained the input vector by concatenation of word embedding and character embedding, entered it into the LSTM-CRF-based NER network to obtain the entity recognition results, and then obtained the relationship between entities through affine attention and softmax. REBEL (Cabot and Navigli, 2021) uses BART as a prototype to transform triplet extraction into a sequence-generating sequence problem through grammar rules. Chargrid (Katti et al., 2018) uses the OCR model of character granularity to extract characters, and CNN of the encoder-decoder is used for text box detection and semantic segmentation of key information for document images. Katti et al. (Katti et al., 2018) proposed using word2vec as a word-level feature extractor and fusing visual information to improve the extraction performance. Bertgrid (Denk and Reisswig, 2019), inspired by Chargrid (Katti et al., 2018), uses the encoder-decoder of a word-granularity convolutional neural network for semantic segmentation. ViBERTgrid (Lin et al., 2021) fuses text features from BERT with image features from convolutional neural networks to achieve better results.
2.2. Multilingual Multi-modal Fusion-based Methods.
The task of form information extraction has developed into a multi-modal task that can integrate CV and NLU (alm; Yang et al., 2024, 2020, 2022b, 2019; Chai et al., 2024), which needs to use the structure information and content information of the form. Most current multimodal tasks on form parsing are extensions of related CV and NLP tasks. The main challenge in form parsing is how to integrate visual language models, Examples such as Vision Transformer (ViT) (Wu et al., 2021) and contrast Language-Image Pretraining CLIP (Radford et al., 2019) have demonstrated the potential to bridge the gap between text and images. These models provide new ways to understand visual elements in the context of textual content.
Figure 2. Overall architecture of the proposed XFormParser. For Multimodal input, XFormParser utilizes layoutXLM to generate vectors containing text, visual, and spatial positional information. Subsequently, these vectors are fed into the downstream joint network to complete SER and RE tasks. The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM and Biaffine.
DocStruct (Wang et al., 2020) integrates visual, layout, and semantic information, utilizing a multilayer perceptron and optical character detection to extract layout and visual information. It uses visual features as mobile features and employs BERT to obtain semantic information. An attention mechanism gate is used to decide whether to use visual features as mobile information. Negative sampling is employed in the training method, achieving hierarchical extraction for form parsing. The effectiveness of key information extraction is validated on the FUNSD (Guillaume Jaume, 2019) dataset. TRIE (Zhang et al., 2020) combines text detection and information extraction, allowing the two tasks to mutually enhance each other for better information extraction results. The StructLM (Li et al., 2021) model learns cell layouts and understands which words belong to the same cell. Its pre-training tasks also learn the positional relationships at the text box level. The LiLT (Wang et al., 2022) model proposes a language-independent transformer focusing on text layout, introducing a new Bidirectional Attention Complement Mechanism (BiACM) to enhance cross-modal cooperation. It also proposes Key Point Localization (KPL) and Cross-modal Alignment Identification (CAI) tasks, combined with the widely used Masked Visual Language Model (MVLM) as pre-training objectives. ERNIE-Layout (Peng et al., 2022), proposed by Baidu, enhances layout knowledge throughout the workflow to learn better representations that combine text, layout, and image features. It reorders the input sequence during serialization and introduces related pre-training tasks, such as reading order prediction, to learn the correct reading order of documents. To improve the model’s layout awareness, spatial disentangled attention is integrated into the multimodal transformer, and the region replacement prediction task is incorporated into the pre-training phase. LayoutLMv3 (Huang et al., 2022) also pre-trains with word-patch alignment objectives, learning cross-modal alignment by predicting whether the corresponding image patches of text words are masked. Its simple unified architecture and training objectives make LayoutLMv3 a general pre-training model suitable for both text-centric and image-centric document AI tasks. XFormParser (Chen et al., 2023) first generates initial relation predictions for entity pairs extracted from document scan images. It then captures global structure knowledge from previous iterative predictions and incorporates this knowledge into entity representations. This ”generate-capture-incorporate” loop is repeated multiple times, allowing entity representations and global structure knowledge to reinforce each other.
3. Method
3.1. Task Definition
The semi-structured form after OCR processing consists of a list of semantic entities, each of which consists of a set of words, the coordinates of the bounding box, and a picture, called the cell of the form and defined as:
(1) $b_{i}=[[w_{1i},...,w_{mi}],[x_{1i},x_{2i},y_{1i},y_{2i}],[image]]$
Where $[w_{1i},\dots,w_{mi}]$ represents the word group, $x_{1i}/x_{2i}$ and $y_{1i}/y_{2i}$ are the upper-left and lower-right coordinates respectively, and image represents the visual information. The documents in the dataset we used are annotated with labels for each entity and relationships between entities. We represent each comment document as:
(2) $D=[[{b_{1},\ldots,b_{n}}],[{l_{1},\ldots,l_{n}}],[{(b_{1},h_{1}),\ldots,(b_{m},b_{h_{m}})}]]$
where $l$ is the label of each entity and $L$ is a predefined set of entity labels. $(bi,bhi)$ refers to the relationship between entities $b_{i}$ and $b_{hi}$, as well as the link point from $b_{hi}$ to $b_{i}$. It is worth noting that the entity may have relationships with multiple entities or may not have a relationship with any other entity.
Cell Classification Task (SER). The cell classification problem needs to classify all the cells and obtain a list of resulting labels $L=[{l_{1},\ldots,l_{n}}]$, where $l_{i}$ corresponds to the cell labels, which can be: Header entity (HEADER), QUESTION entity (QUESTION), ANSWER entity (ANSWER) and OTHER entity (OTHER). When the entity classification is completed, each token is converted into a BIO tag according to the entity category.
Relationship Extraction Task (RE). Given the above definition and description of a form, for a form D, each cell bi needs to find its corresponding sequence of cells in the whole form $B_{b_{i}}=[{b_{j},\ldots,b_{k}}]$, Finally, the normalized cell relation $R_{b_{i}}=[\{(b_{i},b_{j}),\ldots,$
$(b_{i},b_{n})\}]$ is obtained.
3.2. Overall Architecture
As shown in Figure 2, we built the XFormParser model based on the multimodal Transformer architecture. The model accepts information from three different modalities: text, position, and vision, which are encoded using text embedding, 2D position embedding and image embedding layers, respectively. Concatenate the text and image embeddings and then add the position embeddings to obtain the input embeddings. The input embedding is encoded by layoutXLM, and the context representation is output through the dense layer. On the one hand, the representation vector is passed through the MLP layer to obtain the entity classification. On the other hand, the entity embedding vector is mined through the Bi-LSTM sequence relationship and then entered into the $MLP_{head}$ or $MLP_{tail}$ according to the type to obtain the entity expression. The score is obtained by Biaffine (Nguyen and Verspoor, 2019) to determine whether there is a relationship between two entities.
3.3. layoutXLM and Multimodal Input
layoutXLM adds two new embedding layers, 2-D Position Embedding and Image Embedding, based on BERT. The 2-D Position Embedding utilizes OCR and other technologies to obtain the content and position information of each text block in the document. By considering the top-left corner of the document page as the origin of the coordinate system, the representation of each text block in terms of horizontal coordinate, vertical coordinate, width, and height can be calculated, and the final 2-D Position Embedding is the sum of the representations of the four sub-layers. The Image Embedding divides the image into several blocks based on the bounding boxes of each word in the OCR results and corresponds each block to each word. After normalizing embeddings of multiple modalities, they are input into layoutXLM.
Vectorization of Text Information. Vectorization of text information involves converting textual data into numerical representations that can be processed by machine learning algorithms. This transformation is essential for tasks such as text classification, sentiment analysis, and information retrieval. To achieve this, techniques such as tokenization, word embeddings, and sentence embeddings are employed. Tokenization breaks down the text into smaller units like words or subwords.LayoutXLM uses Roberta Tokenizers (Liu et al., 2019) to map these tokens into dense vectors in a continuous vector space where semantically similar words are positioned closely together. Sentence embeddings, on the other hand, provide a vector representation for entire sentences, capturing the semantic meaning of the sentence as a whole. The final vector representation of the text information is a crucial input for various natural language processing models, enabling them to understand and generate human language more effectively.
Vectorization of Location Information. Two-dimensional position embedding uses OCR and other techniques to obtain the content and position information of each text block in the document. The upper-left corner of the document page is regarded as the origin of coordinates to construct the coordinate system, so that the representation of each text block coordinate corresponding to four embedding sub-layers of coordinates, width and height can be calculated, and the final two-dimensional position embedding is the sum of the embedding of four sub-layers.
Vectorization of Image Feature. In this work, the image embedding is not simply obtained by uniform segmentation, but the text block is used to locate the bounding box of each word in the result, and the image is divided into a number of equal size patches, so as to ensure that each word has the embedding vector of the image containing it.
3.4. Form Cell Classifier.
As shown in Figure 2, the pre-trained model receives input [$[\{b_{1},\ldots,$
$b_{n}\}]$, obtains hidden states $H=[h_{1},\ldots,h_{m}]$, and processes the hidden representation H through a fully connected layer (dense). The hidden representation Hner for the NER task is obtained, and then the classification result $logits_{ner}$ is obtained by a classification MLP. $logits_{ner}$ uses $argmax$ to get the predicted $labels=[L_{1},\ldots,L_{n}]$, $L_{i}$, $L_{i}$ corresponds to the labels of each entity, such as ”QUESTION”, ”ANSWER”, etc. In the input layer, to obtain a better entity representation, Leveraging a more expressive pre-trained model, LayoutXLM introduces coordinate information from bounding boxes during pretraining, for this task, we use label embeddings for each entity. First, we get the pre-trained model output of an entity, denoted as $[h^{re}{begin_{i}},\ldots,h^{re}{end_{i}}]$, the model uses a pooling operation to pool the input representations to obtain the entity representation, and then we concatenate the entity representation and the label embedding $l_{i}$. As input to the document encoder for each semantic entity, the formula is as follows,
(3) $e_{i}=pooling([h_{begin_{i}}^{re},...,h_{end_{i}}^{re}])\oplus l_{i}$
3.5. Entity Relation Classification Decoder.
For the relation discrimination task, in order to make the two tasks share H invisible, there should also be a dense layer for the relation discrimination task, $H_{re}=dense_{re}(H)$. $begin_{i}$ represents the index at which the entity starts in $H^{re}$, $end_{i}$ represents the index at which the entity ends in $H$, the vector from $begin_{i}$ to $end_{i}$ in $H^{re}$ is obtained, and the entity expression in the pre-trained model is obtained by pooling (Reimers and Gurevych, 2019). meanpooling is adopted in this paper, where $(l_{i})$ denotes the vector representation obtained by embedding the entity label $(L_{i})$into the label vector space. In the experiment, it was found that the distribution of entity expression and label embedding was not uniform, which would make it difficult for MLP to learn the importance of both information. Therefore, $e_{i}$ was input into BiLSTM as the encoder to unify, and after the entity passed the encoder, According to the type, it enters the head entity encoder $MLP_{head}$ or the tail entity encoder $MLP_{tail}$, and finally obtains the entity representation. It obtains the score through Biaffine (Nguyen and Verspoor, 2019) to determine whether there is a relationship between the two entities.
3.6. Traning Method
Traning Loss. Since it is a joint model, it needs to be trained together to calculate the loss. Cross entropy is used. The loss formula is as follows:
(4) $Loss_{ner}=\text{CE}(logits_{ner},label_{ner})$
(5) $Loss_{re}=\text{CE}(Score_{relations},label_{re})$
(6) $Loss=Loss_{ner}+Loss_{re}$
where CE is the cross-entropy function. $Loss_{ner}$ is the loss for the entity recognition task, $Loss_{re}$ is the loss for the relationship discrimination task, and the final training Loss is the sum of the two used as the learning target of the joint task.
Warm-up soft label.Soft labeling (Huang et al., 2019), also known as soft targets or probabilistic labeling, is a technique in natural language processing where labels for training data are represented as probability distributions rather than as hard, binary labels. Soft labels provide more information about the uncertainty and distribution of classes, leading to better generalization. In traditional hard labels, each sample can only belong to one category, while using soft labels can represent situations where a sample may belong to multiple categories.
We propose a warm-up soft label based on the warming mechanism. In the beginning stage of training, hard labels are used to supervise the model to converge quickly. In the middle and later stages of training, soft labels are used to help model training. During the transition period between soft and hard labels, a warming mechanism is added. , and continuously increase the weight of soft tags.
This is useful for tasks such as multi-label classification and object detection. For the $logits_{ner}$ obtained in the entity classification task, the soft label is calculated and the soft label embedding is calculated. The specific formula is as follows,
(7) $LE_{sl}=\frac{softmax(logits_{ner})\cdot LE_{weight}}{N}$
where $N$ is the number of tags, $LE_{weight}$ is the weight of the word list that stores tag embeddings, $logits_{ner}$ first obtains the distribution probability through a layer of softmax, and $softmax(logits_{ner})$ Perform dot multiplication with $LE_{weight}$ to get a weighted label embedding vector. Finally, the weighted label embedding vector is divided by the number of samples in the dataset $N$ to obtain the average value of the label embedding vector $LE_{softlabel}$, $LE_{sl}$ for short.
In actual experiments, it was found that the training effect was not good in the first half of model training. Although soft label contains more information, it cannot guide the model to learn tasks in the early stage of training, so hard label is still used in the early stage of model training. After the model has been trained to have preliminary capabilities, then use soft label training, and provide a transition for the conversion of hard label and soft label. In fact, the final label embedding calculation method is as follows:
(8) $\alpha=\min(1,(epoch_{now}-epoch_{start}/epoch_{warm}))$
(9) $LE=\begin{cases}LE_{hl},&epoch_{now}\leq epoch_{start}\\ LE_{hl}\times(1-\alpha)+LE_{sl}\times\alpha,&epoch_{now}>epoch_{start}\\ \end{cases}$
where $\alpha$ is a parameter that decays sublinearly with training. where $epoch_{now}$ represents the current training epoch, $epoch_{start}$ denotes the starting epoch, and $epoch_{warm}$ is a predefined value determining the warm-up duration. The parameter $\alpha$ acts as a scaling factor for the learning rate, ensuring a gradual transition from an initial learning rate to the desired rate, as the training progresses.
Table 1. Training configurations for XFormParser
Setup Multi-language model
Optimizer AdamW
Weight ratio 0.1
Lr scheduler LINEAR
vocab size 250002
Max Steps 512
Batch size 8
Initial learning rate 5e-5
Training epochs 100
Evaluation metric F1 score
Table 2. Language-specific fine-tuning F1 accuracy on FUNSD and XFUND (fine-tuning on X, testing on X).“SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
Model FUNSD ZH JA ES FR IT DE PT Avg.
SER XLM-RoBERTaBASE 66.70 87.74 77.61 61.05 67.43 66.87 68.14 68.18 70.47
InfoXLMBASE 68.52 88.68 78.65 62.30 70.15 67.51 70.63 70.08 72.07
LayoutXLMBASE 79.40 89.24 79.21 75.50 79.02 80.82 82.22 79.03 80.56
LiLT[InfoXLM]BASE 84.15 89.38 79.64 79.11 79.53 83.76 82.31 82.20 82.51
GeoLayoutLM 92.86 - - - - - - - -
XFormParser[LiLT] 91.42 91.89 82.25 87.18 87.64 89.42 87.05 87.86 88.09 (+5.58)
XFormParser 92.46 93.14 82.59 87.77 88.69 90.51 88.48 88.68 89.04 (+6.53)
RE XLM-RoBERTaBASE 26.59 51.05 58.00 52.95 49.65 53.05 50.41 39.82 47.69
InfoXLMBASE 29.20 52.14 60.00 55.16 49.13 52.81 52.62 41.70 49.10
LayoutXLMBASE 54.83 70.73 69.63 68.96 63.53 64.15 65.51 57.18 64.32
LiLT[InfoXLM]BASE 62.76 72.97 70.37 71.95 69.65 70.43 65.58 58.74 67.81
GeoLayoutLM 89.45 - - - - - - - -
GOSE[LiLT] 76.97 87.52 80.96 85.95 86.46 84.15 80.23 73.84 82.01
XFormParser[LiLT] 90.02 92.00 91.32 90.21 91.01 91.37 91.11 88.48 90.82 (+8.81)
XFormParser 91.24 93.42 92.19 90.82 91.55 92.48 92.36 89.12 91.65 (+9.64)
Table 3. Multi-Language fine-tuning accuracy (F1) on the XFUND dataset (fine-tuning on 8 languages all, testing on X), where “SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
Model FUNSD ZH JA ES FR IT DE PT Avg.
SER XLM-RoBERTaBASE 66.33 88.3 77.86 62.23 70.35 68.14 71.46 67.26 71.49
InfoXLMBASE 65.38 87.41 78.55 59.79 70.57 68.26 70.55 67.96 71.06
LayoutXLMBASE 79.24 89.73 79.64 77.98 81.73 82.1 83.22 82.41 82.01
LiLT[InfoXLM]BASE 85.74 90.47 80.88 83.40 85.77 87.92 87.69 84.93 85.85
XFormParser 93.89 94.02 90.94 90.19 89.72 91.74 91.94 90.94 91.67 (+5.82)
RE XLM-RoBERTaBASE 36.38 67.97 68.29 68.28 67.27 69.37 68.87 60.82 63.41
InfoXLMBASE 36.99 64.93 64.73 68.28 68.31 66.90 63.84 57.63 61.45
LayoutXLMBASE 66.71 82.41 81.42 81.04 82.21 83.10 78.54 70.44 78.23
LiLT[InfoXLM]BASE 74.07 84.71 83.45 83.35 84.66 84.58 78.78 76.43 81.25
XFormParser 97.00 95.49 94.53 95.67 96.76 97.3 95.49 95.06 95.89 (+14.64)
4. Experiments and Discussion
4.1. Experiment Settings
Datasets
FUNSD (Jaume et al., 2019) (Guillaume Jaume, 2019) is a scanned document dataset for form parsing. It consists of 149 training samples and 50 test samples with various layouts.XFUND (Xu et al., 2021) (Xu et al., 2022) is a multilingual form parsing benchmark. It includes 1,393 fully annotated forms in 7 languages. Each language contains 199 forms, with 149 forms in the training set and 50 forms in the test set.
XFUND and FUNSD datasets have two tasks: Semantic Entity Recognition, where BIO labeling format is used for sequence labeling of each entity, and Relation Extraction, where all possible pairs of given semantic entities are generated to gradually construct a candidate set of relations, identifying all entity pairs with existing associations.
In this paper, the constructed InDFormBench dataset is analyzed, and compared with the scale of the public dataset. We can see that InDFormBench has the following advantages compared with the other two datasets:
1) Larger data size: The training set of the data set of this paper contains 422 samples, and the validation set and the test set contain 70 samples each. In contrast, the training and validation sets of the FUNSD dataset and the XFUND Chinese dataset both have only 150 samples. Larger data sets can provide more samples for the model to learn and train, which helps improve the generalization ability and performance of the model.
2) Chinese language type: Compared with FUNSD data set which is English language type, the data set of this paper is also Chinese language type. For the research or application scenarios of Chinese document processing, the data set of this paper is more in line with the actual needs, because it better reflects the characteristics and language structure of Chinese documents.
3) Complete data set partition: The training set, validation set and test set of the data set in this paper are clearly divided, containing different numbers of samples respectively. This complete dataset partitioning ensures that the model is trained, validated, and tested on different subsets of the data, thus accurately evaluating the performance and robustness of the model.
Table 4. Cross-lingual zero-shot transfer F1 accuracy on FUNSD and XFUND (fine-tuning on FUNSD, testing on X).
Model FUNSD ZH JA ES FR IT DE PT Avg.
SER XLM-RoBERTaBASE 66.70 41.44 30.23 30.55 37.10 27.67 32.86 39.36 38.24
InfoXLMBASE 68.52 44.08 36.03 31.02 40.21 28.80 35.87 45.02 41.19
LayoutXLMBASE 79.40 60.19 47.15 45.65 57.57 48.46 52.52 53.90 55.61
LiLT[InfoXLM]BASE 84.15 61.52 51.84 51.01 59.23 53.71 60.13 63.25 60.61
XFormParser 92.46 72.00 58.25 59.12 69.92 73.72 71.88 73.46 71.35 (+10.74)
RE XLM-RoBERTaBASE 26.59 16.01 26.11 24.40 22.40 23.74 22.88 19.96 22.76
InfoXLMBASE 29.20 24.05 28.51 24.81 24.54 21.93 20.27 20.49 24.23
LayoutXLMBASE 54.83 44.94 44.08 47.08 44.16 40.90 38.20 36.85 43.88
LiLT[InfoXLM]BASE 62.76 47.64 50.81 49.68 52.09 46.97 41.69 42.72 49.30
GOSE[LiLT] 76.97 69.30 68.05 70.72 71.45 63.55 59.97 58.30 67.29
XFormParser 91.24 74.02 81.77 83.02 79.60 78.61 80.18 81.01 81.18 (+9.83)
4.2. Language-specific Fine-tuning
Table 2 presents the results of specific language fine-tuning experiments on the XFUND and FUNSD public datasets. Each column in the table represents a different language (FUNSD, ZH, JA, ES, FR, IT, DE, PT). The table compares different models, including XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, LayoutLMv3BASE, LiLT[InfoXLM]BASE, GeoLayoutLM, and XFormParserBASE. The numerical values in the table indicate the performance metrics of each model fine-tuned on specific languages. XFormParser demonstrates superior performance compared to other models in most languages. On the FUNSD dataset, XFormParser achieves a performance of 92.46%. XFormParser also exhibits strong performance on the XFUND datasets in different languages, with average performance metrics of 89.04% (SER) and 90.54% (RE).
Table 5. Ablation study of our model using LayoutXLM as the backbone on the FUNSD and XFUND (ZH). The symbol EN denotes FUNSD and ZH means Chinese language.
Method Task Component SER F1 Accuracy RE F1 Accuracy↑
SER RE Decoder soft label EN ZH EN ZH
XFormParser 92.46 93.42 91.2 93.14
1 w/o Task RE 91.89 92.86 - -
2 w/o Task SER - - 90.90 92.64
3 w/o Decoder 91.40 92.75 79.79 81.73
3 w/o soft label 91.19 92.19 90.75 91.20
4.3. Multi-Language Fine-tuning
XFUND dataset inspired Multilingual fine-tuning Task, refers to multilingual task fine-tuning on XFUND dataset, all trained on 8 languages and tested on a specific language, where ”SER” stands for semantic entity recognition and ”RE” stands for relation extraction. As shown in Table 3, in the multilingual training task, XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, and XFormParserBASE all show stronger ability than but language fine-tuning, obtaining higher F1 score. In the case of multiple languages, XFormParserBASE shows a more powerful extraction ability. XFormParser model performs best in the multi-language fine-tuning experiment. The average F1 accuracy of SER task is 91.67%. For RE task, XFormParser model also achieves the best performance in multi-language fine-tuning experiments, with an average F1 of 95.89%. These results underscore the efficacy of the XFormParser model in handling multi-language tasks within the XFUND dataset, particularly in comparison to other baseline and advanced models. With more data, our model demonstrates stronger learning capabilities.
4.4. Zero-shot Fine-tuning
Previous experiments illustrate that our method achieves improvements using full training samples. We explored the transferability of our model structure and found that it will gain strong transferability on RE tasks. Thus, we compare with the previous SOTA model LiLT on few-shot settings. The experimental results in Table 4 indicate that the average performance of XFormParser still outperforms the SOTA model LiLT and GOSE. In the task of SER, XFormParser exhibited the highest F1 scores across all languages, achieving an average of 71.35%, with notable improvements in Chinese (72.00%) and Portuguese (73.46%). In the task of RE, the highest performance in the RE task was again seen with XFormParser, which achieved an impressive average F1 score of 81.18%, showing its robustness in relation extraction across different languages. the XFormParser model consistently outperforms other models in both the SER and RE tasks across various languages, highlighting its superior cross-lingual transfer capabilities, and it can improve the generalization of the model.
4.5. Ablation Study
Table 6. For the ablation experiments of epoch start and epoch warm, epoch start refers to the epoch round when soft label is started, epoch warm refers to whether to use the transition mechanism, and the total training rounds are 100. The first line refers to the method of warm up soft label that is not applicable.
epoch start epoch warm RE F1
91.20
10 92.67
20 92.74
30 92.44
30 93.14
40 91.06
40 91.35
50 90.93
50 91.43
To better understand the working principle of the model and determine the extent to which the key components or strategies contribute to the model performance, we designed ablation experiments to verify the feasibility of the method proposed in this paper. In ablation experiments, a series of modifications or eliminations are made to the original model to see how these modifications affect the model’s performance. Ablation experiments are designed and implemented from three innovations of our method:
Effectiveness of individual components. We further investigate the effectiveness of different modules in our method. we compare our model with the following variants in Table 5.
1) w/o multi-task. In this variant, we try to remove the multi-task training method and only use SER or RE for training. This change causes a significant performance decay. The results shown in Table 5 suggest that training two tasks at the same time has the same effect on both SER and RE tasks. With improvement, two tasks that share pre-trained model parameters will learn cross information that improves this task.
2) w/o Decoder. In this variant, we remove the decoder. This change causes a significant performance decay. This suggests the injection of an additional decoder can guide the powerful decoding capabilities of entities and provide strong dependencies for relationship classification.
3) w/o Warm-up soft label. In this variant, we remove the training method Warm-up soft label from XFormParser. This change means the model only uses hard label due to the training. The results shown in Table 5 indicate that Warm-up soft label can improve the effect of the model and prevent the model from overfitting. Experimental data Table 6 indicates that as epoch start increases, model performance initially improves but then decreases. This trend suggests that the model learns sufficient data features by a certain stage, and further training with Warm-up soft label does not enhance performance and may even lead to overfitting.
4.6. Visualization Display
We visualize the SER and RE of XFormParser on a text-intensive form image as shown in Figure 3(b), where the orange boxes are named entities and the arrows represent the matching relationships between the entities. This figure confirms the effectiveness of XFormParser.
Figure 3. Illustration of (a) The form image that is entered into the system; (b) Visualization of SER and RE results.
5. Conclusion
Aiming at the common problems in rich text form parsing tasks, this paper proposes a semi-structured form parser XFormParser based on multi-modal and multi-lingual knowledge. XFormParser integrates layoutXLM pre-trained backbone, cell classifier, and entity relation discrimination decoder, and implements SER and RE tasks for semi-structured form parsing. At the same time, to enrich the experimental data in this field and improve the parsing ability of industrial application scenarios, this paper constructs a Chinese and English multi-scenario form parsing dataset InDFormBench. Three different Settings (such as Language-specific fine-tuning, Multi-Language fine-tuning, and Cross-lingual) are designed on two benchmark datasets and InDFormBench zero-shot), and the results show the effectiveness and superiority of XFormParser.

### v2

@@
 XFormParser: A Simple and Effective Multimodal Multilingual Semi-structured Form Parser
-Abstract.
-In the domain of document AI, semi-structured form parsing plays a crucial role and is employed across a broad spectrum of academic and industrial applications, including structured data management, knowledge retrieval, and content generation.
-This task encompasses a complex form of data mining that leverages techniques from key information extraction (KIE), dealing with inputs that range from plain text to intricate modal data comprising images and structural layouts.
-The advent of pre-trained multimodal models significantly empowers document AI frameworks to adeptly extract pivotal information from form documents presented in diverse formats such as PDFs and images.
-Nonetheless, the endeavor of form parsing is still encumbered by notable challenges like subpar capabilities in multi-lingual parsing and diminished recall in contexts rich in text and visuals.
-In this work, we introduce a simple but effective Multimodal and Multilingual semi-structured FORM PARSER (XFormParser), which anchored on a comprehensive pre-trained language model and innovatively amalgamates semantic entity recognition (SER) and relation extraction (RE) into a unified framework, enhanced by a novel staged warm-up training approach that employs soft labels to significantly refine form parsing accuracy without amplifying inference overhead.
-Furthermore, we have developed a groundbreaking benchmark dataset, named InDFormBench, catering specifically to the parsing requirements of multilingual forms in various industrial contexts.
-Through rigorous testing on established multilingual benchmarks and InDFormBench, XFormParser has demonstrated its unparalleled efficacy, notably surpassing the state-of-the-art (SOTA) models in RE tasks within language-specific setups by achieving an F1 score improvement of up to 1.79%.
-Our framework exhibits exceptionally improved performance across tasks in both multi-language and zero-shot contexts when compared to existing SOTA benchmarks.
-The robustness and reliability of XFormParser are thoroughly validated through comprehensive evaluations on the InDFormBench benchmark.
-[^1] The codes, datasets, and pre-trained models will be released.
-Semi-structured form, Information extraction, Multi-modal, multilingual knowledge, Pre-trained language model
-conference: ; ; submissionid: 286
-1. Introduction
+Abstract
+In the domain of Document AI, parsing semi-structured image form is a crucial Key Information Extraction (KIE) task.
+The advent of pre-trained multimodal models significantly empowers Document AI frameworks to extract key information from form documents in different formats such as PDF, Word, and images.
+Nonetheless, form parsing is still encumbered by notable challenges like subpar capabilities in multilingual parsing and diminished recall in industrial contexts in rich text and rich visuals.
+In this work, we introduce a simple but effective Multimodal and Multilingual semi-structured FORM PARSER (XFormParser), which anchored on a comprehensive Transformer-based pre-trained language model and innovatively amalgamates semantic entity recognition (SER) and relation extraction (RE) into a unified framework.
+Combined with Bi-LSTM, the performance of multilingual parsing is significantly improved.
+Furthermore, we develop InDFormSFT, a pioneering supervised fine-tuning (SFT) industrial dataset that specifically addresses the parsing needs of forms in various industrial contexts.
+XFormParser has demonstrated its unparalleled effectiveness and robustness through rigorous testing on established benchmarks.
+Compared to existing state-of-the-art (SOTA) models, XFormParser notably achieves up to 1.79% F1 score improvement on RE tasks in language-specific settings.
+It also exhibits exceptional cross-task performance improvements in multilingual and zero-shot settings.1
+XFormParser: A Simple and Effective Multimodal Multilingual Semi-structured Form Parser
+Xianfu Cheng1†, Hang Zhang1†, Jian Yang2, Xiang Li2, Weixiao Zhou1, Fei Liu3, Kui Wu1, Xiangyuan Guan1, Tao Sun1, Xianjie Wu1, Tongliang Li4*, Zhoujun Li1,5*, 1CCSE, Beihang University, 2Beihang University, 3Beijing Language and Culture University, 4Beijing Information Science and Technology University, 5Shenzhen Intelligent Strong Technology Co.,Ltd.
+{buaacxf, zhbuaa0, jiaya, xlggg, wxzhou, gxy0615, buaast, wuxianjie, lizj}@buaa.edu.cn, 17861517116@163.com, wukui0099@gmail.com, tonyliangli@bistu.edu.cn
+†Both authors contributed equally to this research.* Corresponding Author.1 The codes, datasets, and pre-trained models are publicly available at https://github.com/zhbuaa0/xformparser.
+1 Introduction
 Document AI is the technology of automatically reading, understanding, and analyzing business documents.
@@
 A key task of Document AI is to parse and extract form information from scanned documents.
-Form parsing is essentially an entity relation mining task that connects the Named Entity Recognition (NER) task (Li et al., 2020) and Key Information Extraction (KIE) task (Cui et al., 2021; Yu et al., 2021; Hong et al., 2022).
+As shown in Figure 1, form parsing is essentially an entity relation mining task that connects the Named Entity Recognition (NER) task Li et al.
+(2020) and KIE task Cui et al.
+(2021); Yu et al.
+(2021); Hong et al.
+(2022).
 Due to the diversity of the layout and the format, poor quality of scanned document images, and complexity of template structures, representing and understanding the unstructured information in documents using generic rules becomes a highly challenging task.
-Figure 1. An illustration of named entity recognition for unstructured forms.
-Unlike traditional NER tasks (Schuster et al., 2013; Esser et al., 2012) which only deal with textual information, and traditional pattern recognition tasks (Dengel and Klein, 2002; Medvet et al., 2011, 2011), form parsing requires consideration of multiple modalities.
+Figure 1: An illustration of named entity recognition for unstructured forms.
+Unlike traditional NER tasks which only deal with textual information, and traditional pattern recognition tasks Medvet et al.
+(2011); Cheng et al.
+(2020), mainstream Document AI methods typically involve using deep neural networks to model elements in documents from the perspectives of computer vision (CV) He et al.
+(2016); Ren et al.
+(2016), natural language processing (NLP) Zhou et al.
+(2023); Zhang et al.
+(2024c), or multimodal fusion Li et al.
+(2024).
 Apart from textual information, the position and layout of text blocks also play a crucial role in semantic interpretation.
-Previously, rules and workflows for each type of document often needed to be hard, requiring updates with changes in specific formats or when dealing with multiple formats.
-Current researchers believe that effective language models for form parsing need to comprehend target entities and adapt to different document formats in contexts involving multiple modalities.
-To address these issues, advanced Form Document AI models and algorithms are expected to automatically classify, extract, and structure information from business documents, minimizing manual intervention.
-Mainstream document intelligence methods typically involve using deep neural networks to model elements in documents from the perspectives of computer vision (He et al., 2016), natural language processing (Devlin et al., 2018), or Multimodal fusion (Atrey et al., 2010).
 Form parsing algorithms based on deep learning initially involved detecting and classifying specific regions of document images.
-Then, it utilized Optical Character Recognition (OCR) models (Li et al., 2023) to extract text information.
+Then, it utilized Optical Character Recognition (OCR) models Li et al.
+(2023); Cheng et al.
+(2024) to extract text information.
 This process is called unstructured to semi-structured, and the resulting form is called semi-structured.
-Subsequently, different categories of form blocks and their corresponding text content were routed to dedicated information extraction modules, constructing a pipeline to process entire-page documents(Xu et al., 2020b; Wang et al., 2022).
-With the development of pre-training techniques (Zoph et al., 2020) and multimodal fusion technologies, researchers have developed a plethora of semi-structured form parsing algorithms based on Transformer (Vaswani et al., 2017).
-(Yu et al., 2021; Zhang et al., 2020) proposed a graph convolutional neural network information extraction model that fuses textual and visual information for form parsing in specific business scenarios.
-(Xu et al., 2020a) introduced the LayoutLM pre-trained model, which combined language models and Transformer, expanding BERT (Devlin et al., 2018) architecture by incorporating layout information to consider spatial relationships between tokens and textual content connections, resulting in outstanding performance in form parsing tasks.
-Furthermore, LayoutLMv2 (Xu et al., 2020b) and LayoutLMv3 (Huang et al., 2022) enhance the integration of multimodal information in the pretraining stage.
-They also add pretraining strategies for text-image alignment and text-image matching, incorporating word token alignment targets.
-LayoutLMv3 is the first multimodal model in Document AI that does not rely on pre-trained CNN (He et al., 2016) or Faster R-CNN (Girshick, 2015) backbones to extract visual features, further promoting the training efficiency and generalizability of the model.
-The XFormParser(Chen et al., 2023) model proposes a Global Structure Knowledge-guided relation extraction framework.
-The loop uses the ”generate-capture-merge” operation so that entity representation and global structural knowledge reinforce each other.
-Although these models have made significant progress in the field of Semi-structured form parsing, as research and applications have become more widespread, most existing methods still face three limitations: 1) The accuracy of key information extraction from complex, multilingual form images remains insufficient; 2) Multimodal models dominated by visual modalities still lag behind text-dominated multimodal models in form parsing tasks in rich-text and long-text scenarios; 3) General Multi-modal Large Language Models (MLLM), like GPT4 (Huang et al., 2023) and their training data are not completely open-source, making it impossible to directly access advanced pre-trained weights and tracks.
-Therefore, based on the self-supervised pre-training model, it is crucial to further mine the representation methods of entities and the relationships between entities in the background of multi-modal and multi-lingual knowledge and to study more effective fine-tuning paradigms in complex scenes for understanding the mechanism of table information extraction and effectively applying it to the industrial field.
-To address these issues, we propose a semi-structured form parser with multimodal and multilingual knowledge, named XFormParser.
-For input data from semi-structured forms, XFormParser utilizes a text-centric pre-trained model to generate vectors containing text information, visual information, and spatial positional information.
-Subsequently, these vectors are fed into the downstream joint network to complete two tasks: Semantic Entity Recognition (SER) and Relation Extraction (RE), namely NER and KIE tasks for semi-structured form, to realize form parsing.
-We use the multilingual document understanding pre-trained model LayoutXLM (Xu et al., 2022) as the backbone of XFormParser.
-The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM (Chung et al., 2014; Sun et al., 2022) and Biaffine (Nguyen and Verspoor, 2019).
-In addition, we further build InDFormBench, a Chinese and English multi-scenario form parsing dataset for industrial applications, based on public benchmarks.
+Subsequently, different categories of form blocks and their corresponding text content were routed to dedicated information extraction modules, constructing a pipeline to process entire-page documents Xu et al.
+(2020b); Wang et al.
+(2022).
+Current researchers believe that effective language models for form parsing must comprehend target entities and adapt to different document formats in contexts involving multiple modalities.
+Advanced Document AI models Wang et al.
+(2020); Zhang et al.
+(2020); Li et al.
+(2021); Peng et al.
+(2022) are expected to automatically classify, extract, and structure information from business documents, minimizing manual intervention.
+Although advanced models have made significant progress in the field of Semi-structured form parsing, as research and applications have become more widespread, most existing methods still face three limitations: 1) The accuracy of key information extraction from complex, multilingual form images remains insufficient; 2) Multimodal models dominated by visual modalities still lag behind text-dominated multimodal models in form parsing tasks in rich-text and long-text scenarios; 3) Multi-modal large language models (MLLM) such as GPT4o Huang et al.
+(2023); Islam and Moushi (2024) and LayoutLLM Fujitake (2024) are difficult to deploy to the end side and achieve fast and high-performance inference via CPUs or low-memory GPUs due to the excessive weight of model parameters.
+Therefore, in industrial applications, it is a crucial research direction to further mine the entity classification and the relations between entities in multimodal and multilingual forms based on simple and effective pre-trained models (PTM) Zoph et al.
+(2020), and to study more effective fine-tuning paradigms in complex scenes.
+To address these issues, we propose a simple but effective semi-structured form parser with multimodal and multilingual knowledge, named XFormParser.
+For input data from semi-structured forms, XFormParser utilizes the multilingual document understanding PTM LayoutXLM Xu et al.
+(2022) to generate vectors containing text, visual, and spatial positional information.
+Subsequently, these vectors are fed into the downstream joint network to complete two tasks: Semantic Entity Recognition (SER) and Relation Extraction (RE), to realize form parsing.
+The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM Sun et al.
+(2022) and Biaffine Nguyen and Verspoor (2019).
+In addition, we further build InDFormSFT, a Chinese and English multi-scenario form parsing SFT dataset for industrial applications, based on public benchmarks.
 Training the model on this dataset helps XFormParser learn semi-structured forms from the real world and achieve new SOTA performance.
 The contributions of this paper are summarized as follows:
-  • We propose XFormParser, an efficient model for multi-language semi-structured form parsing based on a pre-trained Transformer.
-It integrates two tasks: SER and RE, along with the joint loss function and training method of soft labels warm-up in stages.
+  • We propose XFormParser, which integrates two tasks: SER and RE, along with the joint loss function and training method of soft labels warm-up in stages.
 XFormParser effectively enhances form parsing performance without additional inference resources and overhead.
-  • We construct InDFormBench, a cross-scenario form parsing benchmark dataset in both Chinese and English.
-It contains 500 form images collected from eight major industrial application scenarios, along with corresponding annotation information in JSON format semi-automatically generated and rigorously validated using GPT4.
-  • Through experiments and analysis, we confirm that XFormParser achieves the F1 score of at most 1.79% over the SOTA model for RE tasks in Language-specific scenarios.
+  • We construct InDFormSFT, a cross-scenario form parsing SFT dataset in both Chinese and English.
+It contains 562 form images collected from 8 major industrial application scenarios and corresponding annotation information in JSON format that is semi-automatically generated using tools such as GPT4o and rigorously verified by humans.
+  • Through experiments and analysis, we confirm that XFormParser achieves an F1 score of at most 1.79% over the SOTA model for RE tasks in Language-specific scenarios.
 XFormParser achieves significantly better results than SOTA for both dual-task in Multi-language and Zero-shot scenarios.
-We also verify the effectiveness and reliability of XFormParser through experiments on InDFormBench.
-2. Related Work
-2.1. Single Modality-Based Extraction Methods.
-In recent years, various document AI tasks based on natural language understanding (NLU) or computer vision (CV) have made great progress (alm; Yang et al., 2022c, a; Mo et al., 2024; Cheng et al., 2024; Yang et al., 2024; Wang et al., 2023; Ro et al., 2020).
-CloudScan (Palm et al., 2017) attempts to use LSTM to extract information from invoice-type form documents, on this basis, Sage et al.
-(Sage et al., 2019) added contextual semantic and spatial information for information extraction.
-Nguyen et al.
-(Nguyen and Verspoor, 2019), obtained the input vector by concatenation of word embedding and character embedding, entered it into the LSTM-CRF-based NER network to obtain the entity recognition results, and then obtained the relationship between entities through affine attention and softmax.
-REBEL (Cabot and Navigli, 2021) uses BART as a prototype to transform triplet extraction into a sequence-generating sequence problem through grammar rules.
-Chargrid (Katti et al., 2018) uses the OCR model of character granularity to extract characters, and CNN of the encoder-decoder is used for text box detection and semantic segmentation of key information for document images.
-Katti et al.
-(Katti et al., 2018) proposed using word2vec as a word-level feature extractor and fusing visual information to improve the extraction performance.
-Bertgrid (Denk and Reisswig, 2019), inspired by Chargrid (Katti et al., 2018), uses the encoder-decoder of a word-granularity convolutional neural network for semantic segmentation.
-ViBERTgrid (Lin et al., 2021) fuses text features from BERT with image features from convolutional neural networks to achieve better results.
-2.2. Multilingual Multi-modal Fusion-based Methods.
-The task of form information extraction has developed into a multi-modal task that can integrate CV and NLU (alm; Yang et al., 2024, 2020, 2022b, 2019; Chai et al., 2024), which needs to use the structure information and content information of the form.
-Most current multimodal tasks on form parsing are extensions of related CV and NLP tasks.
-The main challenge in form parsing is how to integrate visual language models, Examples such as Vision Transformer (ViT) (Wu et al., 2021) and contrast Language-Image Pretraining CLIP (Radford et al., 2019) have demonstrated the potential to bridge the gap between text and images.
-These models provide new ways to understand visual elements in the context of textual content.
-Figure 2.
-Overall architecture of the proposed XFormParser.
+The ablation experiments on InDFormSFT demonstrate the effectiveness and robustness of XFormParser.
+Figure 2: Overall architecture of the proposed XFormParser.
 For Multimodal input, XFormParser utilizes layoutXLM to generate vectors containing text, visual, and spatial positional information.
@@
 The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM and Biaffine.
-DocStruct (Wang et al., 2020) integrates visual, layout, and semantic information, utilizing a multilayer perceptron and optical character detection to extract layout and visual information.
-It uses visual features as mobile features and employs BERT to obtain semantic information.
-An attention mechanism gate is used to decide whether to use visual features as mobile information.
-Negative sampling is employed in the training method, achieving hierarchical extraction for form parsing.
-The effectiveness of key information extraction is validated on the FUNSD (Guillaume Jaume, 2019) dataset.
-TRIE (Zhang et al., 2020) combines text detection and information extraction, allowing the two tasks to mutually enhance each other for better information extraction results.
-The StructLM (Li et al., 2021) model learns cell layouts and understands which words belong to the same cell.
-Its pre-training tasks also learn the positional relationships at the text box level.
-The LiLT (Wang et al., 2022) model proposes a language-independent transformer focusing on text layout, introducing a new Bidirectional Attention Complement Mechanism (BiACM) to enhance cross-modal cooperation.
-It also proposes Key Point Localization (KPL) and Cross-modal Alignment Identification (CAI) tasks, combined with the widely used Masked Visual Language Model (MVLM) as pre-training objectives.
-ERNIE-Layout (Peng et al., 2022), proposed by Baidu, enhances layout knowledge throughout the workflow to learn better representations that combine text, layout, and image features.
-It reorders the input sequence during serialization and introduces related pre-training tasks, such as reading order prediction, to learn the correct reading order of documents.
-To improve the model’s layout awareness, spatial disentangled attention is integrated into the multimodal transformer, and the region replacement prediction task is incorporated into the pre-training phase.
-LayoutLMv3 (Huang et al., 2022) also pre-trains with word-patch alignment objectives, learning cross-modal alignment by predicting whether the corresponding image patches of text words are masked.
-Its simple unified architecture and training objectives make LayoutLMv3 a general pre-training model suitable for both text-centric and image-centric document AI tasks.
-XFormParser (Chen et al., 2023) first generates initial relation predictions for entity pairs extracted from document scan images.
-It then captures global structure knowledge from previous iterative predictions and incorporates this knowledge into entity representations.
-This ”generate-capture-incorporate” loop is repeated multiple times, allowing entity representations and global structure knowledge to reinforce each other.
-3. Method
-3.1. Task Definition
-The semi-structured form after OCR processing consists of a list of semantic entities, each of which consists of a set of words, the coordinates of the bounding box, and a picture, called the cell of the form and defined as:
-(1) $b_{i}=[[w_{1i},...,w_{mi}],[x_{1i},x_{2i},y_{1i},y_{2i}],[image]]$
-Where $[w_{1i},\dots,w_{mi}]$ represents the word group, $x_{1i}/x_{2i}$ and $y_{1i}/y_{2i}$ are the upper-left and lower-right coordinates respectively, and image represents the visual information.
-The documents in the dataset we used are annotated with labels for each entity and relationships between entities.
-We represent each comment document as:
-(2) $D=[[{b_{1},\ldots,b_{n}}],[{l_{1},\ldots,l_{n}}],[{(b_{1},h_{1}),\ldots,(b_{m},b_{h_{m}})}]]$
-where $l$ is the label of each entity and $L$ is a predefined set of entity labels.
-$(bi,bhi)$ refers to the relationship between entities $b_{i}$ and $b_{hi}$, as well as the link point from $b_{hi}$ to $b_{i}$.
-It is worth noting that the entity may have relationships with multiple entities or may not have a relationship with any other entity.
-Cell Classification Task (SER).
-The cell classification problem needs to classify all the cells and obtain a list of resulting labels $L=[{l_{1},\ldots,l_{n}}]$, where $l_{i}$ corresponds to the cell labels, which can be: Header entity (HEADER), QUESTION entity (QUESTION), ANSWER entity (ANSWER) and OTHER entity (OTHER).
-When the entity classification is completed, each token is converted into a BIO tag according to the entity category.
-Relationship Extraction Task (RE).
-Given the above definition and description of a form, for a form D, each cell bi needs to find its corresponding sequence of cells in the whole form $B_{b_{i}}=[{b_{j},\ldots,b_{k}}]$, Finally, the normalized cell relation $R_{b_{i}}=[\{(b_{i},b_{j}),\ldots,$
-$(b_{i},b_{n})\}]$ is obtained.
-3.2. Overall Architecture
+2 Related Work
+Limited by the lack of training data and the complexity of the corpus, relatively few effective models for parsing multilingual forms have been proposed in the past few years.
+Most applications use a pipeline approach to process form information one modality by one, such as XLM-RoBERTa Conneau et al.
+(2020), which is a multilingual version of RoBERTa Liu et al.
+(2019), and InfoXLM Chi et al.
+(2021), a multilingual pre-trained model Vaswani et al.
+(2017); OpenAI (2023) that maximizes the mutual information between multilingual and multi-granularity texts, which are used as text information processors to analyze forms.
+The Visual-Language models Wu et al.
+(2021); Radford et al.
+(2019); Zhang et al.
+(2024b) demonstrate the potential to effectively align gaps between textual and other modal features, which can be used to integrate form structure information and content information.
+It is further revealed that form information extraction can be achieved by multi-modal techniques combining CV Cheng et al.
+(2022) and NLP Yang et al.
+(2024, 2020, 2022, 2019); Chai et al.
+(2024); Zhang et al.
+(2024a).
+The LiLT Wang et al.
+(2022) model proposes a language-independent transformer focusing on text layout, introducing a new Bidirectional Attention Complement Mechanism to enhance cross-modal cooperation.
+It also proposes Key Point Localization and Cross-modal Alignment Identification tasks, combined with the widely used Masked Visual Language Model as pre-training objectives.
+LayoutLM Xu et al.
+(2020a) introduced the PTM, which combined language models and Transformer, expanding BERT Devlin et al.
+(2018) architecture by incorporating layout information to consider spatial relations between tokens and textual content connections, resulting in outstanding performance in form parsing tasks.
+Furthermore, LayoutLMv2 Xu et al.
+(2020b) and LayoutLMv3 Huang et al.
+(2022) enhance the integration of multimodal information in the pretraining stage.
+They also add pretraining strategies for text-image alignment and text-image matching, incorporating word token alignment targets.
+Regarding the joint modeling of SER and RE, existing work primarily targets sequence labeling tasks, training jointly at the token granularity Jiang et al.
+(2024); Ji et al.
+(2024); Wang et al.
+(2023); El Khbir et al.
+(2024); Liu et al.
+(2023).
+At present, there is a lack of available models or methods for joint pattern training with cell information.
+GeoLayoutLM Luo et al.
+(2023) improves the feature representation of text and layout by explicitly modeling geometric relationships and special pre-training tasks, which can improve the performance of form information extraction.
+However, the complex model structure makes the pre-training cost of related tasks high and too dependent on data.
+GOSE Chen et al.
+(2023) first generates initial relation predictions for entity pairs extracted from document scan images.
+It then captures global structure knowledge from previous iterative predictions and incorporates it into entity representations.
+This "generate-capture-incorporate" loop is repeated multiple times, allowing entity representations and global structure knowledge to reinforce each other.
+3 Method
+3.1 Task Definition
+The semi-structured form processed by the OCR model can be represented as a list of $n$ semantic entities.
+Each entity consists of a set of words named $Words_{i}$, the coordinates of the bounding box named $Bbox_{i}$, and an image named $Image_{i}$, called $i$-th cell of the form and defined as $b_{i}$:
+$b_{i}=[Words_{i},Bbox_{i},Image_{i}]$ (1)
+The documents in our dataset are annotated with labels for each entity and relations between entities. We represent each comment document $D$ as:
+$D=[B,L,R]$ (2)
+where $B$=$[b_{1},\ldots,b_{n}]$ denote all the cells of the form; $L$=$[{l_{1},\ldots,l_{n}}]$ is a predefined set of entity labels and $l$ is the label of each entity; $R$=$[{(b_{1},b_{2}),\ldots,(b_{j},b_{k})}]$ is the set of relations between entities, $(b_{j},b_{k})$ refers to the relation between $j$-th entity and $k$-th entity, as well as the link point from $b_{k}$ to $b_{j}$.
+It is worth noting that one entity may have relations with multiple entities or may not have a relationship with some other entity.
+Semantic Entity Recognition Task.
+The semantic entity recognition needs to classify all the cells and obtain a list of resulting labels $L$, the label of cells can be: Header entity (HEADER), QUESTION entity (QUESTION), ANSWER entity (ANSWER) and OTHER entity (OTHER).
+When the entity classification is completed, each token needs to be converted into a BIO label according to the entity category, to facilitate the alignment with experiments on open-source benchmark datasets.
+Relation Extraction Task.
+Given the above definition and description of a form, for a form $D$, each cell $b_{i}$ needs to find its corresponding sequence of cells in the whole form $B_{b_{i}}$=$[{b_{j},\ldots,b_{k}}]$, Finally, the normalized cell relation $R_{b_{i}}$=[$[\{(b_{i},b_{j}),\ldots,(b_{i},b_{k})\}]$ is obtained.
+3.2 Overall Architecture
 As shown in Figure 2, we built the XFormParser model based on the multimodal Transformer architecture.
-The model accepts information from three different modalities: text, position, and vision, which are encoded using text embedding, 2D position embedding and image embedding layers, respectively.
+The model accepts information from three different modalities: text, position, and vision, encoded using text embedding, 2D position embedding, and image embedding layers, respectively.
 Concatenate the text and image embeddings and then add the position embeddings to obtain the input embeddings.
@@
 On the one hand, the representation vector is passed through the MLP layer to obtain the entity classification.
-On the other hand, the entity embedding vector is mined through the Bi-LSTM sequence relationship and then entered into the $MLP_{head}$ or $MLP_{tail}$ according to the type to obtain the entity expression.
-The score is obtained by Biaffine (Nguyen and Verspoor, 2019) to determine whether there is a relationship between two entities.
-3.3. layoutXLM and Multimodal Input
+On the other hand, the entity embedding vector is mined through the Bi-LSTM sequence relation and then entered into the $MLP_{head}$ or $MLP_{tail}$ according to the type to obtain the entity expression.
+The $Score$ is obtained by Biaffine to determine whether there is a relation between two entities.
+3.3 PTM and Multimodal Input
 layoutXLM adds two new embedding layers, 2-D Position Embedding and Image Embedding, based on BERT.
-The 2-D Position Embedding utilizes OCR and other technologies to obtain the content and position information of each text block in the document.
+The 2-D Position Embedding corresponds to text blocks with the content and position information in the document, both content and position information obtained by OCR and other technologies.
 By considering the top-left corner of the document page as the origin of the coordinate system, the representation of each text block in terms of horizontal coordinate, vertical coordinate, width, and height can be calculated, and the final 2-D Position Embedding is the sum of the representations of the four sub-layers.
-The Image Embedding divides the image into several blocks based on the bounding boxes of each word in the OCR results and corresponds each block to each word.
+The image embedding divides the image into several blocks based on the bounding boxes of each word in the OCR results, and each block corresponds to each word.
 After normalizing embeddings of multiple modalities, they are input into layoutXLM.
 Vectorization of Text Information.
-Vectorization of text information involves converting textual data into numerical representations that can be processed by machine learning algorithms.
-This transformation is essential for tasks such as text classification, sentiment analysis, and information retrieval.
-To achieve this, techniques such as tokenization, word embeddings, and sentence embeddings are employed.
-Tokenization breaks down the text into smaller units like words or subwords.LayoutXLM uses Roberta Tokenizers (Liu et al., 2019) to map these tokens into dense vectors in a continuous vector space where semantically similar words are positioned closely together.
-Sentence embeddings, on the other hand, provide a vector representation for entire sentences, capturing the semantic meaning of the sentence as a whole.
-The final vector representation of the text information is a crucial input for various natural language processing models, enabling them to understand and generate human language more effectively.
+Tokenization, word embeddings, and sentence embeddings mainly realize the vectorization of text information.
+Among them, tokenization breaks down text into smaller units such as words or subwords.
+LayoutXLM uses RoBERTa Tokenizers to map these tokens into dense vectors in a continuous vector space where semantically similar words are positioned closely together.
+Sentence embeddings are also used, provide a vector representation for entire sentences, capturing the semantic meaning of the sentence as a whole.
 Vectorization of Location Information.
 Two-dimensional position embedding uses OCR and other techniques to obtain the content and position information of each text block in the document.
-The upper-left corner of the document page is regarded as the origin of coordinates to construct the coordinate system, so that the representation of each text block coordinate corresponding to four embedding sub-layers of coordinates, width and height can be calculated, and the final two-dimensional position embedding is the sum of the embedding of four sub-layers.
+The upper-left corner of the document page is regarded as the origin of constructing the coordinate system so that the embeddings of each text block corresponding to XY coordinates, width, and height can be calculated, and the final two-dimensional position embedding is the sum of the four embeddings.
 Vectorization of Image Feature.
-In this work, the image embedding is not simply obtained by uniform segmentation, but the text block is used to locate the bounding box of each word in the result, and the image is divided into a number of equal size patches, so as to ensure that each word has the embedding vector of the image containing it.
-3.4. Form Cell Classifier.
-As shown in Figure 2, the pre-trained model receives input [$[\{b_{1},\ldots,$
-$b_{n}\}]$, obtains hidden states $H=[h_{1},\ldots,h_{m}]$, and processes the hidden representation H through a fully connected layer (dense).
-The hidden representation Hner for the NER task is obtained, and then the classification result $logits_{ner}$ is obtained by a classification MLP.
-$logits_{ner}$ uses $argmax$ to get the predicted $labels=[L_{1},\ldots,L_{n}]$, $L_{i}$, $L_{i}$ corresponds to the labels of each entity, such as ”QUESTION”, ”ANSWER”, etc.
-In the input layer, to obtain a better entity representation, Leveraging a more expressive pre-trained model, LayoutXLM introduces coordinate information from bounding boxes during pretraining, for this task, we use label embeddings for each entity.
-First, we get the pre-trained model output of an entity, denoted as $[h^{re}{begin_{i}},\ldots,h^{re}{end_{i}}]$, the model uses a pooling operation to pool the input representations to obtain the entity representation, and then we concatenate the entity representation and the label embedding $l_{i}$.
-As input to the document encoder for each semantic entity, the formula is as follows,
-(3) $e_{i}=pooling([h_{begin_{i}}^{re},...,h_{end_{i}}^{re}])\oplus l_{i}$
-3.5. Entity Relation Classification Decoder.
-For the relation discrimination task, in order to make the two tasks share H invisible, there should also be a dense layer for the relation discrimination task, $H_{re}=dense_{re}(H)$.
-$begin_{i}$ represents the index at which the entity starts in $H^{re}$, $end_{i}$ represents the index at which the entity ends in $H$, the vector from $begin_{i}$ to $end_{i}$ in $H^{re}$ is obtained, and the entity expression in the pre-trained model is obtained by pooling (Reimers and Gurevych, 2019).
-meanpooling is adopted in this paper, where $(l_{i})$ denotes the vector representation obtained by embedding the entity label $(L_{i})$into the label vector space.
+In this work, the image embedding is not simply obtained by uniform segmentation, but the text block is used to locate the bounding box of each word in the result, and the image is divided into several equal size patches, to ensure that each word has embedding vector of the image containing it.
+3.4 Semantic Entity Recognizer
+As shown in Figure 2, the PTM receives multi-modal input $B$=$[{b_{1},\ldots,b_{n}}]$, obtains hidden states $H$, and processes the hidden representation $H$ through a fully connected layer (Dense).
+The hidden representation $H_{ser}$ for the SER task is obtained, and then the classification result $logits_{ser}$ is obtained by a classification MLP.
+The formulas are expressed as follows:
+$H=PLM(B)$ (3)
+$H_{ser}=Dense_{ser}(H)$ (4)
+$logits_{ser}=MLP_{ser}(H_{ser})$ (5)
+Then $logits_{ser}$ uses $argmax$ to get the predicted vectors $P$=$[p_{1},\ldots,p_{n}]$, $p_{i}$ corresponds to the label $l_{i}$ of $i$-th entity, such as "QUESTION", "ANSWER", etc.
+3.5 Relation Extraction Decoder
+For the RE task, we first get the PTM output of the $i$-th entity, denoted as $H_{re}$.
+The model then uses a pooling operation Reimers and Gurevych (2019) on $H_{re}$ to obtain the embeddings of the entities.
+Then we concatenate the entity embeddings and the corresponding label embeddings $p_{i}$ predicted by the SER.
+As the output of the encoder for each semantic entity, the formula is as follows:
+$e_{i}=pooling(H_{re})\oplus p_{i}$ (6)
+Note that for two tasks to implicitly share $H$, $H_{re}$ needs to be obtained using a Dense layer identical to Eq.
+(4), that is, $H_{re}$= $Dense_{re}(H)$.
+In addition, mean-pooling is experimentally verified to be the most effective pooling operation applied in Eq.
+(6).
 In the experiment, it was found that the distribution of entity expression and label embedding was not uniform, which would make it difficult for MLP to learn the importance of both information.
-Therefore, $e_{i}$ was input into BiLSTM as the encoder to unify, and after the entity passed the encoder, According to the type, it enters the head entity encoder $MLP_{head}$ or the tail entity encoder $MLP_{tail}$, and finally obtains the entity representation.
-It obtains the score through Biaffine (Nguyen and Verspoor, 2019) to determine whether there is a relationship between the two entities.
-3.6. Traning Method
-Traning Loss. Since it is a joint model, it needs to be trained together to calculate the loss. Cross entropy is used. The loss formula is as follows:
-(4) $Loss_{ner}=\text{CE}(logits_{ner},label_{ner})$
-(5) $Loss_{re}=\text{CE}(Score_{relations},label_{re})$
-(6) $Loss=Loss_{ner}+Loss_{re}$
+Therefore, $e_{i}$ was input into the Bi-LSTM decoder to unify, and after the entity passed the decoder, according to the type, it entered the head entity decoder $MLP_{head}$ or the tail entity decoder $MLP_{tail}$, and finally produces the entity representation.
+It obtains the score through Biaffine to determine whether there is a relation between the two entities.
+3.6 Traning Method
+Training Loss. Since it is a joint model, the team must be trained to calculate the loss. The loss formula is as follows:
+$Loss_{ser}=\text{CE}(logits_{ser},L_{ser})$ (7)
+$Loss_{re}=\text{CE}(Score_{Biaffine},L_{re})$ (8)
+$Loss=Loss_{ser}+Loss_{re}$ (9)
 where CE is the cross-entropy function.
-$Loss_{ner}$ is the loss for the entity recognition task, $Loss_{re}$ is the loss for the relationship discrimination task, and the final training Loss is the sum of the two used as the learning target of the joint task.
-Warm-up soft label.Soft labeling (Huang et al., 2019), also known as soft targets or probabilistic labeling, is a technique in natural language processing where labels for training data are represented as probability distributions rather than as hard, binary labels.
-Soft labels provide more information about the uncertainty and distribution of classes, leading to better generalization.
-In traditional hard labels, each sample can only belong to one category, while using soft labels can represent situations where a sample may belong to multiple categories.
-We propose a warm-up soft label based on the warming mechanism.
-In the beginning stage of training, hard labels are used to supervise the model to converge quickly.
+$Loss_{ser}$ is the loss for the SER task, $Loss_{re}$ is the loss for the RE task, and the final training Loss is the sum of the two used as the learning target of the joint task.
+Warm-up Soft Label.
+We propose an improved warm-up soft label Huang et al.
+(2019) based on the warming mechanism, which can effectively improve the performance when applied to fine-tuning form parsing models.
+In the beginning stage of training, hard labels are used to supervise the model so that it can converge quickly.
 In the middle and later stages of training, soft labels are used to help model training.
-During the transition period between soft and hard labels, a warming mechanism is added.
-, and continuously increase the weight of soft tags.
-This is useful for tasks such as multi-label classification and object detection.
-For the $logits_{ner}$ obtained in the entity classification task, the soft label is calculated and the soft label embedding is calculated.
-The specific formula is as follows,
-(7) $LE_{sl}=\frac{softmax(logits_{ner})\cdot LE_{weight}}{N}$
-where $N$ is the number of tags, $LE_{weight}$ is the weight of the word list that stores tag embeddings, $logits_{ner}$ first obtains the distribution probability through a layer of softmax, and $softmax(logits_{ner})$ Perform dot multiplication with $LE_{weight}$ to get a weighted label embedding vector.
-Finally, the weighted label embedding vector is divided by the number of samples in the dataset $N$ to obtain the average value of the label embedding vector $LE_{softlabel}$, $LE_{sl}$ for short.
+During the transition period between soft and hard labels, a warming mechanism is added, and the weight of soft tags is continuously increased.
+This is useful for tasks such as multi-label classification.
 In actual experiments, it was found that the training effect was not good in the first half of model training.
-Although soft label contains more information, it cannot guide the model to learn tasks in the early stage of training, so hard label is still used in the early stage of model training.
+Although soft labels contain more information, they cannot guide the model in learning tasks in the early stage of training, so hard labels are still used in the early stage of model training.
 After the model has been trained to have preliminary capabilities, then use soft label training, and provide a transition for the conversion of hard label and soft label.
-In fact, the final label embedding calculation method is as follows:
-(8) $\alpha=\min(1,(epoch_{now}-epoch_{start}/epoch_{warm}))$
-(9) $LE=\begin{cases}LE_{hl},&epoch_{now}\leq epoch_{start}\\ LE_{hl}\times(1-\alpha)+LE_{sl}\times\alpha,&epoch_{now}>epoch_{start}\\ \end{cases}$
-where $\alpha$ is a parameter that decays sublinearly with training.
-where $epoch_{now}$ represents the current training epoch, $epoch_{start}$ denotes the starting epoch, and $epoch_{warm}$ is a predefined value determining the warm-up duration.
-The parameter $\alpha$ acts as a scaling factor for the learning rate, ensuring a gradual transition from an initial learning rate to the desired rate, as the training progresses.
-Table 1. Training configurations for XFormParser
+For the construction method of the warm-up soft label and related training parameter Settings, please refer to Appendix A.
 Setup Multi-language model
@@
 Training epochs 100
-Evaluation metric F1 score
-Table 2. Language-specific fine-tuning F1 accuracy on FUNSD and XFUND (fine-tuning on X, testing on X).“SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
+Evaluation metric ref. B.4
+Table 1: Training configurations for XFormParser
 Model FUNSD ZH JA ES FR IT DE PT Avg.
@@
 XFormParser 91.24 93.42 92.19 90.82 91.55 92.48 92.36 89.12 91.65 (+9.64)
-Table 3.
-Multi-Language fine-tuning accuracy (F1) on the XFUND dataset (fine-tuning on 8 languages all, testing on X), where “SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
+Table 2: Language-specific fine-tuning F1 accuracy on FUNSD and XFUND (fine-tuning on X, testing on X).“SER” denotes the semantic entity recognition, and “RE” denotes the relation extraction.
 Model FUNSD ZH JA ES FR IT DE PT Avg.
@@
 XFormParser 97.00 95.49 94.53 95.67 96.76 97.3 95.49 95.06 95.89 (+14.64)
-4. Experiments and Discussion
-4.1. Experiment Settings
-Datasets
-FUNSD (Jaume et al., 2019) (Guillaume Jaume, 2019) is a scanned document dataset for form parsing.
-It consists of 149 training samples and 50 test samples with various layouts.XFUND (Xu et al., 2021) (Xu et al., 2022) is a multilingual form parsing benchmark.
+Table 3: Multi-language fine-tuning accuracy (F1) on the XFUND dataset (fine-tuning on 8 languages all, testing on X), where “SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
+4 Experiments and Discussion
+4.1 Experiment Settings
+Datasets.
+FUNSD Guillaume Jaume (2019) is a scanned document dataset for form parsing.
+It is a subset of the RVL-CDIP Harley et al.
+(2015) and consists of 149 training samples and 50 test samples with various layouts.
+XFUND Xu et al.
+(2022) is a multilingual form parsing benchmark.
 It includes 1,393 fully annotated forms in 7 languages.
 Each language contains 199 forms, with 149 forms in the training set and 50 forms in the test set.
-XFUND and FUNSD datasets have two tasks: Semantic Entity Recognition, where BIO labeling format is used for sequence labeling of each entity, and Relation Extraction, where all possible pairs of given semantic entities are generated to gradually construct a candidate set of relations, identifying all entity pairs with existing associations.
-In this paper, the constructed InDFormBench dataset is analyzed, and compared with the scale of the public dataset.
-We can see that InDFormBench has the following advantages compared with the other two datasets:
-1) Larger data size: The training set of the data set of this paper contains 422 samples, and the validation set and the test set contain 70 samples each.
-In contrast, the training and validation sets of the FUNSD dataset and the XFUND Chinese dataset both have only 150 samples.
-Larger data sets can provide more samples for the model to learn and train, which helps improve the generalization ability and performance of the model.
-2) Chinese language type: Compared with FUNSD data set which is English language type, the data set of this paper is also Chinese language type.
-For the research or application scenarios of Chinese document processing, the data set of this paper is more in line with the actual needs, because it better reflects the characteristics and language structure of Chinese documents.
-3) Complete data set partition: The training set, validation set and test set of the data set in this paper are clearly divided, containing different numbers of samples respectively.
-This complete dataset partitioning ensures that the model is trained, validated, and tested on different subsets of the data, thus accurately evaluating the performance and robustness of the model.
-Table 4. Cross-lingual zero-shot transfer F1 accuracy on FUNSD and XFUND (fine-tuning on FUNSD, testing on X).
+XFUND and FUNSD datasets have two tasks: SER, where BIO labeling format is used for sequence labeling of each entity, and RE, where all possible pairs of given semantic entities are generated to gradually construct a candidate set of relations, identifying all entity pairs with existing associations.
+In this paper, InDFormSFT is constructed based on the Chinese-English context and eight application scenarios, the training set contains 422 samples, and the validation set and test set contain 70 samples each.
+Compared with FUNSD and XFUND datasets, larger datasets can provide more samples for the model to learn and train, which helps to improve the generalization ability and performance of the model.
+Please refer to Appendix B for the construction process of InDFormSFT.
 Model FUNSD ZH JA ES FR IT DE PT Avg.
@@
 GOSE[LiLT] 76.97 69.30 68.05 70.72 71.45 63.55 59.97 58.30 67.29
-XFormParser 91.24 74.02 81.77 83.02 79.60 78.61 80.18 81.01 81.18 (+9.83)
-4.2. Language-specific Fine-tuning
+XFormParser 91.24 74.02 81.77 83.02 79.60 78.61 80.18 81.01 81.18 (+13.89)
+Table 4: Cross-lingual zero-shot transfer F1 accuracy on FUNSD and XFUND (fine-tuning on FUNSD, testing on XFUND).
+4.2 Language-specific Fine-tuning
 Table 2 presents the results of specific language fine-tuning experiments on the XFUND and FUNSD public datasets.
-Each column in the table represents a different language (FUNSD, ZH, JA, ES, FR, IT, DE, PT).
+Each column in the table represents a different language (FUNSD, ZH, JA, ES, FR, IT, DE, PT) Xu et al.
+(2022).
 The table compares different models, including XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, LayoutLMv3BASE, LiLT[InfoXLM]BASE, GeoLayoutLM, and XFormParserBASE.
-The numerical values in the table indicate the performance metrics of each model fine-tuned on specific languages.
+The numerical values in the table indicate the performance metrics of each model fine-tuned in specific languages.
 XFormParser demonstrates superior performance compared to other models in most languages.
@@
 XFormParser also exhibits strong performance on the XFUND datasets in different languages, with average performance metrics of 89.04% (SER) and 90.54% (RE).
-Table 5. Ablation study of our model using LayoutXLM as the backbone on the FUNSD and XFUND (ZH). The symbol EN denotes FUNSD and ZH means Chinese language.
 Method Task Component SER F1 Accuracy RE F1 Accuracy↑
@@
 4 w/o soft label 91.19 92.19 90.75 91.20
-4.3. Multi-Language Fine-tuning
-XFUND dataset inspired Multilingual fine-tuning Task, refers to multilingual task fine-tuning on XFUND dataset, all trained on 8 languages and tested on a specific language, where ”SER” stands for semantic entity recognition and ”RE” stands for relation extraction.
-As shown in Table 3, in the multilingual training task, XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, and XFormParserBASE all show stronger ability than but language fine-tuning, obtaining higher F1 score.
+Table 5: Ablation study of our model using LayoutXLM as the backbone, InDFormSFT as training dataset and test on the FUNSD and XFUND (ZH). The symbol EN denotes FUNSD and ZH means Chinese language.
+4.3 Multi-language Fine-tuning
+XFUND dataset inspired Multilingual fine-tuning Task, refers to multilingual task fine-tuning on XFUND dataset, all trained on 8 languages and tested on a specific language.
+As shown in Table 3, in the multilingual training task, XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, and XFormParserBASE all show stronger ability than language-specific fine-tuning, obtaining higher F1 score.
 In the case of multiple languages, XFormParserBASE shows a more powerful extraction ability.
-XFormParser model performs best in the multi-language fine-tuning experiment.
+XFormParser performs best in the multi-language fine-tuning experiment.
 The average F1 accuracy of SER task is 91.67%.
-For RE task, XFormParser model also achieves the best performance in multi-language fine-tuning experiments, with an average F1 of 95.89%.
-These results underscore the efficacy of the XFormParser model in handling multi-language tasks within the XFUND dataset, particularly in comparison to other baseline and advanced models.
+For RE task, XFormParser also achieves the best performance in multi-language fine-tuning experiments, with an average F1 of 95.89%.
+These results underscore the efficacy of the XFormParser in handling multi-language tasks within the XFUND dataset, particularly in comparison to other baseline and advanced models.
 With more data, our model demonstrates stronger learning capabilities.
-4.4. Zero-shot Fine-tuning
+4.4 Zero-shot Fine-tuning
 Previous experiments illustrate that our method achieves improvements using full training samples.
@@
 In the task of RE, the highest performance in the RE task was again seen with XFormParser, which achieved an impressive average F1 score of 81.18%, showing its robustness in relation extraction across different languages.
-the XFormParser model consistently outperforms other models in both the SER and RE tasks across various languages, highlighting its superior cross-lingual transfer capabilities, and it can improve the generalization of the model.
-4.5. Ablation Study
-Table 6.
-For the ablation experiments of epoch start and epoch warm, epoch start refers to the epoch round when soft label is started, epoch warm refers to whether to use the transition mechanism, and the total training rounds are 100.
-The first line refers to the method of warm up soft label that is not applicable.
+the XFormParser consistently outperforms other models in both the SER and RE tasks across various languages, highlighting its superior cross-lingual transfer capabilities, and it can improve the generalization of the model.
+4.5 Ablation Study
 epoch start epoch warm RE F1
@@
 50 91.43
+Table 6: For the ablation experiments of epoch start and epoch warm, epoch start refers to the epoch round when the soft label is started, epoch warm refers to whether to use the transition mechanism, and the total training rounds are 100.
+The first line refers to the method of warm-up soft label that is not applicable.
 To better understand the working principle of the model and determine the extent to which the key components or strategies contribute to the model performance, we designed ablation experiments to verify the feasibility of the method proposed in this paper.
@@
 The results shown in Table 5 suggest that training two tasks at the same time has the same effect on both SER and RE tasks.
-With improvement, two tasks that share pre-trained model parameters will learn cross information that improves this task.
+With improvement, two tasks that share PTM parameters will learn cross-information that improves this task.
 2) w/o Decoder.
@@
 This change causes a significant performance decay.
-This suggests the injection of an additional decoder can guide the powerful decoding capabilities of entities and provide strong dependencies for relationship classification.
+This suggests the injection of an additional decoder can guide the powerful decoding capabilities of entities and provide strong dependencies for relation classification.
 3) w/o Warm-up soft label.
 In this variant, we remove the training method Warm-up soft label from XFormParser.
-This change means the model only uses hard label due to the training.
-The results shown in Table 5 indicate that Warm-up soft label can improve the effect of the model and prevent the model from overfitting.
-Experimental data Table 6 indicates that as epoch start increases, model performance initially improves but then decreases.
-This trend suggests that the model learns sufficient data features by a certain stage, and further training with Warm-up soft label does not enhance performance and may even lead to overfitting.
-4.6. Visualization Display
-We visualize the SER and RE of XFormParser on a text-intensive form image as shown in Figure 3(b), where the orange boxes are named entities and the arrows represent the matching relationships between the entities.
+This change means the model only uses hard labels due to the training.
+The results shown in Table 5 indicate that a Warm-up soft label can improve the effect of the model and prevent the model from overfitting.
+Experimental data Table 6 indicates that as epoch starts increases, model performance initially improves but then decreases.
+This trend suggests that the model learns sufficient data features by a certain stage, and further training with a Warm-up soft label does not enhance performance and may even lead to overfitting.
+4.6 Visualization Display
+We visualize the SER and RE of XFormParser on a text-intensive form image as shown in Figure 3(b), where the orange boxes are named entities and the arrows represent the matching relations between the entities.
 This figure confirms the effectiveness of XFormParser.
-Figure 3. Illustration of (a) The form image that is entered into the system; (b) Visualization of SER and RE results.
-5. Conclusion
+Figure 3: Illustration of (a) The form image that is entered into the system; (b) Visualization of SER and RE results.
+5 Conclusion
 Aiming at the common problems in rich text form parsing tasks, this paper proposes a semi-structured form parser XFormParser based on multi-modal and multi-lingual knowledge.
-XFormParser integrates layoutXLM pre-trained backbone, cell classifier, and entity relation discrimination decoder, and implements SER and RE tasks for semi-structured form parsing.
-At the same time, to enrich the experimental data in this field and improve the parsing ability of industrial application scenarios, this paper constructs a Chinese and English multi-scenario form parsing dataset InDFormBench.
-Three different Settings (such as Language-specific fine-tuning, Multi-Language fine-tuning, and Cross-lingual) are designed on two benchmark datasets and InDFormBench zero-shot), and the results show the effectiveness and superiority of XFormParser.
+XFormParser integrates layoutXLM pre-trained backbone, semantic entity recognizer, and relation extraction decoder, and implements SER and RE tasks for semi-structured form parsing.
+At the same time, to enrich the experimental data in this field and improve the parsing ability of industrial application scenarios, this paper constructs a Chinese and English multi-scenario form parsing SFT dataset InDFormSFT.
+Four different Settings (such as Language-specific fine-tuning, Multi-language fine-tuning, Cross-lingual fine-tuning, and Zero-shot) are designed on two benchmark datasets and InDFormSFT ), and the results show the effectiveness and superiority of XFormParser.
+6 Limitations
+The purpose of this work is to provide a simple, efficient, and easy-to-deploy semi-structured form parsing component for the end side (PC or mobile).
+Although we use a multi-modal approach and expand the training set to improve the performance of the model, while taking into account the multi-language parsing scenario, this work still has the following limitations.
+Diversity of Languages.
+InDFormSFT only includes two languages, Chinese and English, and lacks expansion of form knowledge for the other six languages in XFUND.
+After that, multi-language augmented data can be constructed by using the same data set construction process with the help of machine translation and layout design algorithms.
+Model Diversity.
+The comparative experiments do not include the experimental results of the large model of multi-modal documents on relevant benchmarks, thus lacking the most powerful demonstration of the upper bound of the performance of the model for the current task.
+In addition, our work does not further improve the multilingual pre-trained model backbone and directly adopts the strongest and most easy-to-use model investigated as the backbone of XFormParser.
+These limitations need to be further studied and improved.
+Completeness of Verification.
+There is a lack of validation of model compression and inference acceleration methods such as model distillation, model pruning, and model quantization.
+Acknowledgments
+This work was partially supported by the National Natural Science Foundation of China (Grant Nos.
+62276017, 62406033, U1636211, 61672081, 62272025), and the State Key Laboratory of Complex& Critical Software Environment (Grant No.
+SKLCCSE-2024ZX-18).
+Appendix A Build the Warm-up Soft Label
+Soft labeling, also known as soft targets or probabilistic labeling, is a technique in natural language processing where labels for training data are represented as probability distributions rather than as hard, binary labels.
+Soft labels provide more information about the uncertainty and distribution of classes, leading to better generalization.
+In traditional hard labels, each sample can only belong to one category, while using soft labels can represent situations where a sample may belong to multiple categories.
+We used the improved warm-up soft label, which can effectively improve the performance when applied to fine-tuning form parsing models.
+In the beginning stage of training, hard labels are used to supervise the model so that it can converge quickly.
+In the middle and later stages of training, soft labels are used to help model training.
+During the transition period between soft and hard labels, a warming mechanism is added, and the weight of soft tags is continuously increased.
+This is useful for tasks such as multi-label classification.
+For the soft labels $logits_{ser}$ and its embedding are calculated.
+The specific formula is as follows,
+$LE_{sl}=\frac{softmax(logits_{ser})\cdot LE_{weight}}{N}$ (10)
+where $N$ is the number of tags, $LE_{weight}$ is the weight of the word list that stores tag embeddings, $logits_{ser}$ first obtains the distribution probability through a layer of softmax, and $softmax(logits_{ser})$ Perform dot multiplication with $LE_{weight}$ to get a weighted label embedding vector.
+Finally, the weighted label embedding vector is divided by the number of samples in the dataset $N$ to obtain the average value of the label embedding vector $LE_{softlabel}$, $LE_{sl}$ for short.
+Figure 4: It shows the process of data search.
+On the basis of the constructed data, the title of the document is extracted as the search term, and other form files are searched in the document search engine.
+Figure 5: Firstly, the optical character recognition tool is used to process the file, and the text information and border information are obtained.
+The data structure of the form is constructed through the border and text, including the structure information and the text information of the form.
+Figure 6: It shows the construction of a form data auxiliary labeling tool.
+Firstly, the optical character recognition tool is used to process the file to obtain text and border information.
+Then the form data filtering tool is used to determine whether the cell can be labeled.
+Then the form information extraction model is used for auxiliary labeling, and finally, the data result is obtained by manual labeling.
+Figure 7: Firstly, a PDF form is obtained, the text and border information are extracted through OCR, and the cell characteristics are calculated to determine whether it conforms to the form structure.
+If it does not meet the conditions, it is directly discarded.
+If it meets the conditions, the vacancy rate of the cell is calculated.
+The qualified forms are pre-predicted and labeled by the model, and the pre-labeled results are imported into the labeling system.
+Finally, the correct labeled data are obtained by manual verification for training and testing.
+In actual experiments, it was found that the training effect was not good in the first half of model training.
+Although soft labels contain more information, they cannot guide the model in learning tasks in the early stage of training, so hard labels are still used in the early stage of model training.
+After the model has been trained to have preliminary capabilities, then use soft label training, and provide a transition for the conversion of hard label and soft label.
+In fact, the final label embedding calculation method is as follows:
+$\alpha=\min(1,(ep-ep_{start}/ep_{warm}))$ (11)
+$LE=\begin{cases}LE_{hl},&ep\leq ep_{start}\\ \alpha LE_{sl}+\beta LE_{hl},&ep>ep_{start}\\ \end{cases}$ (12)
+where $\alpha$ is a parameter that decays sublinearly with training and $\beta$=$1$-$\alpha$.
+where $ep$ represents the current training epoch, $ep_{start}$ denotes the starting epoch, and $ep_{warm}$ is a predefined value determining the warm-up duration.
+The parameter $\alpha$ acts as a scaling factor for the learning rate, ensuring a gradual transition from an initial learning rate to the desired rate, as the training progresses.
+Appendix B Approach to building InDFormSFT
+The large labeled data set is the main support for the high performance of deep learning.
+Form datasets are a common type of datasets used to collect, store, and analyze user-submitted data.
+It is commonly used in various application areas such as market research, user surveys, online registration, order forms, etc.
+In recent years, there have been some academic data sets in the field of forms.
+This kind of data set contains two kinds of information, one is the image of the original document, and the other is the specific information of the document that has been annotated or parsed, usually including the image, the coordinates of the text box, the text content of the text box, the label of the text box and the relationship between the text box.
+B.1 Data Collection and Annotation
+The basic data set is constructed through the Chinese and English data sets of FUNSD and XFUND, and then the Chinese form data on the Internet is collected.
+To avoid privacy and sensitive information issues of real-world documents, this paper collects documents publicly available on the Internet.
+Semi-structured forms were collected through Baidu Library, and the search keywords included: a comprehensive table of university teachers, a comprehensive table of senior safety engineers, a comprehensive table of professor-level senior engineers, etc.
+By downloading and collecting the files of the Shenzhen Stock Exchange, Shanghai Stock Exchange, and other financial platforms, the form files that meet the task definition are found.
+Figures 456 show the data search process, the form data filtering, and the construction process of the auxiliary labeling tool, respectively.
+Finally, to facilitate the pre-processing and labeling of forms, a set of engineering developments of data screening and labeling process based on Chinese forms was completed, as shown in Figure 7 below.
+B.2 Instances of Semi-structured Data
+The form data also aligns with the format of the XFUND dataset, with cell granularity, where each cell contains information such as absolute cell coordinates (box), cell text information (text), cell label information (label), cell ids (id), and linking between cells (linking).
+The first row of the picture is shown in Figure 8(a), and the corresponding annotated data format is shown in Figure 8(b).
+Figure 8: Illustration of (a) A form image in InDFormSFT; (b) The annotation corresponding to the first row of the image.
+Dataset partition Data volume Question Entity Answer Entity Single Entity Title Entity Continuous char entity
+Training set 422 6702 6825 422 370 194
+Validation set 70 1375 1443 65 18 83
+Testing set 70 1468 1644 78 58 18
+Table 7: Analysis of the number of entity labels in InDFormSFT.
+Dataset partition One-to-one One-to-two One-to-three One-to-many
+Training set 11806 265 96 171
+Validation set 2366 55 38 39
+Testing set 2626 76 53 72
+Table 8: Entity relationship analysis of InDFormSFT.
+In Figure 8(b), lang denotes the language of the text.
+In this case, its value is "zh", indicating that the text is Chinese.
+version indicates the version of the dataset, where the value is "0.1" and "split", which means the dataset is split into train, val, and test sets.
+"id" represents a unique identifier for the form data.
+documents contain a list of form cell information, and box represents the coordinates of the text’s bounding box on the image.
+The value is a list of four integers representing the top-left x-coordinate, the top-left y-coordinate, the bottom-right x-coordinate, and the bottom-right y-coordinate.
+text represents the text content of the cell.
+Here, the value is a string that contains some text.
+label denotes the label content of the cell, which includes the SINGLE entity (SINGLE), QUESTION entity (QUESTION), ANSWER entity (ANSWER), and continuous character entity (ANSWERNUM) mentioned above.
+id represents the unique identifier of the cell.
+linking represents a linking relationship between cells and is a list of entity pairs with two ids, the first for the question entity and the second for the answer or consecutive character entity.
+img contains the form image information, where fname represents the name of the image file, width represents the width of the image, and height represents the height of the image.
+B.3 Analysis of InDFormSFT
+As shown in Table 7, for the analysis of the entity content of the data set, according to the table content analysis, the entity labels of the data set of this paper have the following characteristics: the number of question entities and answer entities is large.
+In the training set, the number of question entities is 6702, the number of answer entities is 6825, and the number of question entities and answer entities is basically the same.
+This indicates that there are a large number of question-and-answer entities in the dataset that need to be identified and annotated.
+These entities may include person names, place names, organizations, etc., and we need our model to be able to accurately identify and label these entities.
+The number of single entities, title entities, and continuous character entities is relatively rare, and recognizing these sparse entities is a challenging task.
+According to the content analysis of Table 8, the table shows the division of the data set and the corresponding relationship types.
+Entities are classified by relationship type, including one-to-one, one-to-two, one-to-three and one-to-many (greater than three).
+These relation types describe the number of entity-time correspondences in the form information extraction task.
+We can see that most relationships are concentrated in one-to-one relationships, and a few exist in one-to-two, one-to-three, and one-to-many (greater than three) relationships.
+By analyzing the distribution of the number of different correspondences, we can get the distribution of the number of samples of different relation types in the dataset.
+B.4 Evaluation Metrics
+In this paper, the target tasks for the datasets used are divided into SER and RE.
+For the SER task, the model needs to determine the class of each Cell in the form: SINGLE, QUESTION, ANSWER, and ANSWERNUM.
+The evaluation metric is Cell Acc (CA).
+The accuracy of Cell Discrimination is determined by Correct Cell Discrimination (CCD) and Total Cell Count (TCC).
+The formula is as follows:
+$CA=\frac{CCD}{TCC}$ (13)
+For the cell relation linking task, we need to determine which combinations of cells in each table have a key-value relation.
+The F1-Score is a commonly used metric to evaluate the performance of classification models, which takes into account both Precision and Recall.
+The formula for calculating the F1 score is as follows:
+$F1=2\times\frac{Precision\times Recall}{Precision+Recall}$ (14)
+$Precision=\frac{TP}{TP+FP}$ (15)
+$Recall=\frac{TP}{TP+FN}$ (16)
+In this task, TP represents the number of entity pairs correctly predicted by the model as having a relationship.
+Actually, having a relationship, FP represents the number of entity pairs predicted by the model as having a relationship but actually having no relationship.
+FN represents the number of entity pairs predicted by the model as having no relationship but actually having a relationship.
+The F1 value ranges from 0 to 1, with values closer to 1 indicating better performance of the model.